PERTURBATION AND PREDISSOCIATION OF THE Na_{2} 3^{3}Π_{g} AND 4^{3}Σ_{g}^{+} STATES

Y. LIU, J. LI, D. CHEN, and \underline{LI LI}, Department of Physics and Center for Atomic and Molecular Sciences, Tsinghua University, Beijing 100084, China; R. W. FIELD, Department of Chemistry, MIT, Cambridge Mass 02139; and A. M. LYYRA, Department of Physics, Temple University, Philadelphia PA 19122.

The Na_{2} 3^{3}Π_{g} and 4^{3}Σ_{g}^{+} states dissociate adiabatically to the 3s+4p atomic limit. Energy levels below the 3s+3d atomic limit of these two states have been observed by perturbation facilitated optical-optical double resonance (PFOODR) fluorescence excitation spectroscopy.^{[}^{a}^{,}^{b}^{]}. Energy levels above the 3s+3d atomic limit could not be observed by detecting molecular fluorescence but have been observed recently by detecting atomic 3d → 3p fluorescence. The 3^{3}Π_{g} and 4^{3}Σ_{g}^{+} states interact with the 3s+3d 2^{3}Π_{g} and 3^{3}Σ_{g}^{+} states, respectively. As a result of the perturbation the 3^{3}Π_{g} state strongly predissociates and the linewidth of levels close to the 3s+3d limit is about 15cm^{−1}. The 4^{3}Σ_{g}^{+} state is weakly predissociated.

aLi Li and R. W. Field, J. Mol. Spectrosc. \underline{117}, 245 (1986).

bLi Li and M. Li, J. Mol. Spectrosc. \underline{173}, 25 (1996).